# Low-Complexity Super-Resolution Signature Estimation of XL-MIMO FMCW Radar

### v1

@@ Abstract
I Introduction
Modern radar systems, particularly in automotive applications and emerging sensing technologies, demand ultra-high resolution for reliable environment perception and object detection [1]. This has given rise to a high level of interest in extra-large multiple-input multiple-output (XL-MIMO) radar systems, which leverage hundreds or thousands of antenna elements to achieve fine angular and range resolution while maintaining wide coverage [2]. Enhancing radar resolution necessitates deploying ultra-large arrays combined with high modulation bandwidths [3, 4, 5, 6, 7, 8, 9, 10]. However, this combination introduces two major challenges: (i) the spatial wideband effect (SWE) where the spatial delay across the radar aperture becomes comparable to the range resolution [11], and (ii) the near field effect (NFE) where, at large array scales, the planar wavefront assumption fails for nearby targets [12]. In particular, the SWE significantly complicates joint range and angle-of-arrival (AoA) estimation since the range and angle terms of the intermediate frequency (IF) signal become cross-coupled, making conventional signal processing approaches ineffective. In narrowband MIMO-FMCW radars, joint range-angle estimation is commonly achieved using 2D frequency estimation or sequential 1D techniques with manageable complexity [13]. Conversely, under SWE in wideband radars, the decoupling assumption fails, and the radio scene’s sparsity structure evolves into a block sparsity model [11, 3, 4, 7].
@@ I Introduction
To the best of our knowledge, no existing work in the literature addresses low-complexity, joint range-angle estimation in XL-MIMO FMCW radars under SWE without assuming the number of targets. Motivated by this gap, we first develop a comprehensive system model capturing SWE effects in XL-MIMO FMCW radar systems, that takes into account the inherent range-angle coupling in far-field scenarios. To address the associated high-dimensional estimation challenges, a scalable and computationally efficient algorithm is developed for the joint estimation of the number of targets, their ranges, AoAs, and complex reflection coefficients without requiring prior target count information. This is accomplished by extending the conventional two-dimensional orthogonal matching pursuit (2D-OMP) technique to exploit the block-sparse structure of spatial wideband radar scenes, enabling robust super-resolution signature estimation. Extensive simulation results are presented to validate the proposed method, demonstrating superior estimation accuracy and computational efficiency compared to existing subspace-based and wideband AoA estimation techniques across various XL-MIMO radar configurations.
@@ II System Model
[…]
@@
×$\;\;\times\exp{\left(j2\pi f_{c}\tau^{k}_{q}\right)},$
@@
We, henceforth, deal with only discrete-time signals and use $n$ to denote the discrete-time index with $10\leq n\leq N-$ 1 where $N=\lceil f_{s}T_{ch}\rceil$. Moreover, throughout this work, only stationary targets are considered, and as a result, the Doppler-induced delay is neglected. Since the Doppler-induced delay is affected by SWE, further investigations are required, which we leave for future research. Next, we examine terms I, II, and III.
@@
$=\exp{\left(j2\pi\Omega_{R}^{k}n\right)}\exp{\left(j2\pi\frac{\Omega^{k}_{\theta}}{f_{c}}\gamma\frac{n}{f_{s}}\right)}$
$=\exp{\left(j2\pi\Omega_{R}^{k}n\right)}\exp{\left(j2\pi\Omega^{k}_{\theta}\frac{\alpha}{N}qn\right)}.$
@@
$\exp{\left(-j2\pi\gamma\frac{2R_{k}qd\sin\theta_{k}}{c^{2}}\right)}.$
Here $\exp{(-j\pi\gamma{4R_{k}^{2}}/{c^{2}})}$ does not vary with antenna index and hence can be absorbed in $\widetilde{a}_{k}$. Further, the maximum value of the remaining two terms in the exponents is close to zero and hence can be ignored (This can be verified numerically for example by letting $R_{k}=20m$, $\theta_{k}=90^{\circ}$ $Q=128$, and $\gamma=100MHz/\mu s$).
@@
$\exp{\left(j2\pi f_{c}\tau^{k}_{q}\right)}=\exp{\left(j2\pi f_{c}\tau^{R}_{k}\right)}\exp{\left(j2\pi\frac{qd\sin(\theta_{k})}{\lambda}\right)}.$
@@
×$\;\;\times\underbrace{\exp{(j2\pi\frac{\alpha}{N}\Omega^{k}_{\theta}qn)}}_{SW\;Term}+w_{q}[n],$ (3)
@@ III Proposed Methodology
III-A DFT-based Coarse estimation
@@
$\underbrace{\exp{(j2\pi\frac{\alpha}{N}\Omega_{k}^{\theta}qn)}}_{SW\;Term}+w_{q}[n]\Big{)}\;\times\exp{\left(-j\frac{2\pi}{Q}uq\right)}\exp{\left(-j\frac{2\pi}{N}vn\right)}$ (5)
@@
The number of target-induced blocks in the range-angle response can be estimated using techniques such as clustering [21], support detection [22], or custom peak-finding algorithms [19]. In this work, we assume that no two target blocks overlap under the SWE condition, ensuring distinct separation for detection. To identify these blocks, we employ the 2D peak search method proposed in [23] to locate the peaks of $||\mathbf{Y}_{k}||_{2}^{2}$ within the range-angle response, thereby estimating the number of targets $\hat{K}$. The peak location for each target block provides a coarse estimate of its corresponding range bin ($\hat{u}_{k}$) and AoA bin ($\hat{v}_{k}$), in the neighborhood $\mathcal{N}_{k}$ corresponding to the $k$-th target’s return.
@@
III-B CS-based fine-tuned estimation
@@
For this, we first initialize $\check{y}_{q}[n]=y_{q}[n]$, then for the $k^{th}$ target, we compensate for the SWE using the coarse angle estimate $\hat{\Omega}_{\theta}^{k}$ as
$\tilde{y}_{q}[n]=$ $\check{y}_{q}[n]\times\exp{\left(-j2\pi\frac{\alpha}{N}\hat{\Omega}_{k}^{\theta}qn\right)}$ (6)
@@
×$\times\exp{\left(j2\pi\frac{\alpha}{N}\Omega_{j}^{\theta}qn\right)}$ $\exp{\left(-j2\pi\frac{\alpha}{N}\hat{\Omega}_{\theta}^{k}qn\right)}+\tilde{w}_{q}[n]$
$=\sum_{j=0}^{K-1}\widetilde{a}^{*}_{k}\exp{(j2\pi\Omega^{k}_{R}t)}$ $\exp{(j2\pi\Omega^{k}_{\theta}q)}\underbrace{\exp{\left(j2\pi\frac{\alpha}{N}(\Omega_{\theta}^{j}-\hat{\Omega}_{\theta}^{k})qn\right)}}_{Residue\;Term}$
$+\tilde{w}_{q}[n]$ (7)
For the close values of coarse angle estimates ($\Omega_{\theta}^{j}\approx\Omega_{\theta}^{k}$), the residue term in (7) becomes close to unity and hence the SWE is compensated for the $k^{th}$ path. It should be noted here that the power for the $k^{th}$ target is enhanced significantly compared to other targets due to near-perfect SWE compensation, and the IF signal corresponding to this target can be approximately recovered by putting an appropriate threshold, i.e., update the $\tilde{y}_{q}[n]$
$\tilde{y}_{q}[n]\leftarrow\begin{cases}\tilde{y}_{q}[n],&\text{if }|\tilde{y}_{q}[n]|\geq\gamma\\ 0,&\text{if }|\tilde{y}_{q}[n]|<\gamma\end{cases}$
Equivalently, the IF signal for the $k^{th}$ path, converted to the narrowband signal model, can be written as
@@
This process can be illustrated with an example, as shown in Fig. 2. We consider two targets, where Target 1 is located at [2.02m, $[2.02\,\mathrm{m},\,30^{\degree}]$ and Target 2 at $[0.97\,\mathrm{m},\,-45^{\degree}]$. The system bandwidth is set to $0.1f_{c}$, with $f_{c}=77$ GHz, and the number of fast-time and spatial samples are $R,N=128$. The two unit-amplitude targets produce spread responses in the angle-delay domain due to the Spatial Wideband Effect, as depicted in Fig. 2(a). The coarse angle estimates for Target 1 and Target 2 are $31.03^{\degree}$ and $-47.25^{\degree}$, respectively. Spatial wideband compensation, performed using (7), is shown in Fig. 2(b) and Fig. 2(c). It can be clearly observed that the power concentration of the respective target in the angle-delay domain significantly improves, and the equivalent narrowband target model, as represented by (8), is effectively achieved.
It is important to note that the accuracy of the wideband effect compensation relies heavily on the precision of the coarse angle estimate. At higher angles and larger bandwidths, the spread caused by SWE becomes sufficiently large to induce notable coarse angle bin migration. As a result, the rotation-based array processing technique proposed in [11] becomes ineffective under these conditions, which we demonstrate numerically in Section IV.
@@
where, $\mathbf{Y}^{w}$ is the stacked measurement matrix of the spatial wideband signal in (3), $\mathbf{\Theta}(\Omega_{\theta})\doteq exp(j2\pi\frac{\alpha}{N}\Omega_{\theta}qn)$ is the wideband phase matrix, $\mathbf{b,c}$ are steering vectors as defined in (10), and $\circ$ denotes the element-wise product.
After obtaining the refined signature of the $k^{\text{th}}$ path, $\{\hat{\hat{a}}^{k},\hat{\hat{\Omega}}_{R}^{k},\hat{\hat{\Omega}}_{\theta}^{k}\}$, we remove the corresponding contribution from the intermediate frequency (IF) signal. This operation ensures that the associated peak is eliminated, allowing the algorithm to detect remaining paths in subsequent iterations. Consequently, we update $\check{y}_{q}[n]$ as follows
@@
IV Results
The normalized bandwidth parameter is set to $\alpha=0.1$, and the carrier frequency is fixed at $f_{c}=77\mathrm{GHz}$. A virtual array consisting of $Q=128$ elements is considered. Additionally, the number of fast-time samples per chirp is set to $N=128$, unless otherwise specified.
@@ IV Results
We present an example of localizing $K=2$ targets in Fig. 3. In the spatial narrowband case (Fig. 3(a)), the 2D-MUSIC, 2D-Rotation, and the proposed method all accurately localize the two targets in the range-angle domain. However, under the spatial wideband scenario, the narrowband 2D-MUSIC algorithm detects only one target, while the 2D-Rotation method exhibits a localization error of approximately $4^{\circ}$ for the second target, as illustrated in Fig. 3(b). In contrast, the proposed method successfully localizes both targets with high accuracy, placing them in close proximity to their actual positions in the spatial wideband case.
Figure 3: Target localization at (a) $\alpha=0.01$ (b) $\alpha=0.1$.
Figure 4: RMSE of the estimated parameters at $\alpha=0.1$ (a) Range (b) AOA (c) Complex Coefficient.
Figure 5: (a) Hit rate and (b) false rate at $\alpha=0.1$.
@@
We evaluate the effectiveness of our proposed method by measuring its hit rate, false rate, and Root Mean Square Error (RMSE). The target scenario consists of $K=3$ far-field point targets with ranges uniformly drawn from the interval $[1\,\mathrm{m},20\,\mathrm{m}]$ and angles of arrival (AoAs) uniformly drawn from $[0^{\circ},80^{\circ}]$. The target complex gains are modeled as $a_{k}=\exp{\left(j\psi_{k}\right)}$, where $\psi_{k}\sim\mathcal{U}[0,2\pi)$, while the noise samples are drawn as $w_{q}[t]\sim\mathcal{CN}(\mathbf{0},\sigma^{2}\mathbf{I})$, independently across all virtual array elements and time samples.
We evaluate the performance of the proposed compressive sensing (CS)-based estimation method for XL-MIMO FMCW radar under the Spatial Wideband Effect (SWE) by comparing it with three benchmark techniques: the rotation-based spatial wideband estimation method from [11], the 2D-MUSIC algorithm [27], and the conventional 2D-OMP method [28]. The comparison is conducted in terms of target detection probability and parameter estimation accuracy.
A detected target is classified as a hit if its estimated range and AoA fall within the predefined resolution limits ($0.02:m$ and $1^{\circ}$ in this experiment); otherwise, it is considered a false alarm. The hit rate and false rate for different methods are presented in Fig. 5, while the corresponding RMSE values, computed for correctly detected targets, are shown in Fig. 4.
As shown in Fig. 4(a), the proposed method achieves substantially lower range RMSE across the entire SNR range, with rapid error reduction beyond 00 dB, while 2D-MUSIC and 2D-OMP exhibit consistently higher errors due to their inability to handle the SWE-induced coupling. The 2D-Rotation method improves upon these narrowband techniques but remains inferior to the proposed approach.
@@
Fig. 5 illustrates the target detection performance in terms of hit rate and false rate versus SNR for $\alpha=0.1$. As shown in Fig. 5(a), the proposed method achieves a hit rate close to 1 at SNRs above $-10$ dB, with a rapid performance improvement as SNR increases. In contrast, the 2D-Rotation method maintains a moderate hit rate of approximately 50–60% across the entire SNR range, while both 2D-OMP and 2D-MUSIC consistently fail to reliably detect targets under spatial wideband conditions, achieving hit rates below 20%.
In terms of false detection, depicted in Fig. 5(b), the proposed method exhibits a sharp reduction in false rate with increasing SNR, achieving near-zero false alarms beyond 0 dB. The 2D-Rotation method, however, retains a high false rate of around 40–50%, while 2D-OMP and 2D-MUSIC show persistently high false rates (around 80–85%) across all SNR values. These results clearly demonstrate the superiority of the proposed compressive sensing-based estimation technique for SWE-affected XL-MIMO FMCW radar, delivering both high detection reliability and low false alarm rates, even under challenging spatial wideband conditions.
@@
Using the 2D-MUSIC algorithm for 2D range-angle parameter estimation with $Q=128$, $N=128$, and $K=3$ takes approximately 22 seconds on a 10th-generation Intel i7 processor. In comparison, the 2D-Rotation method performs the same task in just 0.17 seconds. Both of these methods, however, are limited to spatial narrowband scenarios. The proposed method not only reduces the estimation time to 0.15 seconds but also extends applicability to spatial wideband cases, offering both improved efficiency and broader utility.
V Conclusion
In this work, we have proposed a low-complexity joint range-angle estimation methodology for XL-MIMO FMCW radars under the SWE, capable of simultaneously estimating the number of targets and their complex reflection coefficients. Numerical results demonstrate that the proposed technique achieves a significantly higher hit rate and a lower false alarm rate along with lower MSE compared to existing methods. Moreover, the scalability and computational efficiency of the proposed approach make it well-suited for a range of localization-based applications across both low and high bandwidth settings. Additionally, the proposed algorithm is computationally efficient, as it does not require any supplementary mechanisms to mitigate range-angle bin migration effects induced by SWE.

### v2

@@ Abstract
1 Introduction
Modern radar systems, particularly in automotive applications and emerging sensing technologies, demand ultra-high resolution for reliable environment perception and object detection [1]. This has given rise to a high level of interest in extra-large multiple-input multiple-output (XL-MIMO) radar systems, which leverage hundreds or thousands of antenna elements to achieve fine angular and range resolution while maintaining wide coverage [2]. Enhancing radar resolution necessitates deploying ultra-large arrays combined with high modulation bandwidths [3, 4, 5, 6, 7, 8, 9, 10]. However, this combination introduces two major challenges: (i) the spatial wideband effect (SWE) where the spatial delay across the radar aperture becomes comparable to the range resolution [11], and (ii) the near field effect (NFE) where, at large array scales, the planar wavefront assumption fails for nearby targets [12]. While the NFE becomes increasingly important for extremely large arrays and short-range targets, its explicit modeling and joint treatment with SWE are beyond the scope of this work and will be investigated in future studies. On the other hand, the SWE significantly complicates joint range and angle-of-arrival (AoA) estimation since the range and angle terms of the intermediate frequency (IF) signal become cross-coupled, making conventional signal processing approaches ineffective. In narrowband MIMO-FMCW radars, joint range-angle estimation is commonly achieved using 2D frequency estimation or sequential 1D techniques with manageable complexity [13]. Conversely, under SWE in wideband radars, the decoupling assumption fails, and the radio scene’s sparsity structure evolves into a block sparsity model [11, 3, 4, 7].
@@ 1 Introduction
To the best of our knowledge, no existing work in the literature addresses low-complexity, joint range-angle estimation in XL-MIMO FMCW radars under SWE without assuming the number of targets. Motivated by this gap, we first develop a comprehensive system model capturing SWE effects in XL-MIMO FMCW radar systems, which takes into account the inherent range-angle coupling in far-field scenarios. To address the associated high-dimensional estimation challenges, a scalable and computationally efficient algorithm is developed for the joint estimation of the number of targets, their ranges, AoAs, and complex reflection coefficients without requiring prior target count information. This is accomplished by extending the conventional two-dimensional orthogonal matching pursuit (2D-OMP) technique to exploit the block-sparse structure of spatial wideband radar scenes, enabling robust super-resolution signature estimation. Extensive simulation results are presented to validate the proposed method, demonstrating superior estimation accuracy and computational efficiency compared to existing subspace-based and wideband AoA estimation techniques across various XL-MIMO radar configurations.
@@ 2 System Model
$s(t)=\exp{\left(j2\pi\left(f_{c}t+\frac{\gamma}{2}t^{2}\right)\right)},\;\;0\leq t\leq T_{ch},$ (1)
@@
$\;\;\times\exp{\left(j2\pi f_{c}\tau^{k}_{q}\right)},$ (5)
@@
We, henceforth, deal with only discrete-time signals and use $n$ to denote the discrete-time index with $0\leq n\leq N-1$ where $N=\lceil f_{s}T_{ch}\rceil$. Moreover, throughout this work, only stationary targets are considered, and as a result, the Doppler-induced delay is neglected. Since the Doppler-induced delay is affected by SWE, further investigations are required, which we leave for future research. Next, we examine terms I, II, and III.
@@
$=\exp{\left(j2\pi\Omega_{R}^{k}n\right)}\exp{\left(j2\pi\frac{\Omega^{k}_{\theta}}{f_{c}}\gamma\frac{qn}{f_{s}}\right)}$
$=\exp{\left(j2\pi\Omega_{R}^{k}n\right)}\exp{\left(j2\pi\Omega^{k}_{\theta}\frac{\alpha}{N}qn\right)}.$ (7)
@@
$\exp{\left(-j2\pi\gamma\frac{2R_{k}qd\sin\theta_{k}}{c^{2}}\right)}.$ (8)
Here, $\exp{(-j\pi\gamma{4R_{k}^{2}}/{c^{2}})}$ does not vary with antenna index and hence can be absorbed in $\widetilde{a}_{k}$. Further, the maximum value of the remaining two terms in the exponents is close to zero and hence can be ignored (This can be verified numerically, for example by letting $R_{k}=20m$, $\theta_{k}=90^{\circ}$, $Q=128$, and $\gamma=100MHz/\mu s$).
@@
$\exp{\left(j2\pi f_{c}\tau^{k}_{q}\right)}=\exp{\left(j2\pi f_{c}\tau^{R}_{k}\right)}\exp{\left(j2\pi\frac{qd\sin(\theta_{k})}{\lambda}\right)}.$ (9)
@@
$\;\;\times\underbrace{\exp{(j2\pi\frac{\alpha}{N}\Omega^{k}_{\theta}qn)}}_{SW\;Term}+w_{q}[n],$ (10)
@@ 3 Proposed Methodology
3.1 DFT-based coarse estimation
@@
$\underbrace{\exp{(j2\pi\frac{\alpha}{N}\Omega_{k}^{\theta}qn)}}_{SW\;Term}+w_{q}[n]\Big)\;\times\exp{\left(-j\frac{2\pi}{Q}uq\right)}\exp{\left(-j\frac{2\pi}{N}vn\right)},$ (12)
@@
The number of target-induced blocks in the range-angle response can be estimated using techniques such as clustering [21], support detection [22], or custom peak-finding algorithms [19]. In this work, we assume that no two target blocks overlap under the SWE condition, ensuring distinct separation for detection. To identify these blocks, we employ the 2D peak search method proposed in [23] to locate the peaks of $\mathbf{Y}_{k}||_{2}^{2}$ within the range-angle response, thereby estimating the number of targets $\hat{K}$. The peak location for each target block provides a coarse estimate of its corresponding range bin ($\hat{u}_{k}$) and AoA bin ($\hat{v}_{k}$), in the neighborhood $\mathcal{N}_{k}$ corresponding to the $k$-th target’s return.
@@
3.2 CS-based fine-tuned estimation
@@
For this, we first initialize $\check{y}_{q}[n]=y_{q}[n]$, then for the $k$-th target, we compensate for the SWE using the coarse angle estimate $\hat{\Omega}_{\theta}^{k}$ as
$\tilde{y}_{q}[n]=$ $\check{y}_{q}[n]\times\exp{\left(-j2\pi\frac{\alpha}{N}\hat{\Omega}^{k}_{\theta}qn\right)}$ (13)
@@
$\times\exp{\left(j2\pi\frac{\alpha}{N}\Omega_{j}^{\theta}qn\right)}$ $\exp{\left(-j2\pi\frac{\alpha}{N}\hat{\Omega}_{\theta}^{k}qn\right)}+\tilde{w}_{q}[n]$
$=\sum_{j=0}^{K-1}\widetilde{a}^{*}_{k}\exp{(j2\pi\Omega^{j}_{R}t)}$ $\exp{(j2\pi\Omega^{j}_{\theta}q)}\underbrace{\exp{\left(j2\pi\frac{\alpha}{N}(\Omega_{\theta}^{j}-\hat{\Omega}_{\theta}^{k})qn\right)}}_{Residue\;Term}$
$+\tilde{w}_{q}[n]$ . (14)
For the close values of coarse angle estimates ($\Omega_{\theta}^{j}\approx\hat{\Omega}_{\theta}^{k}$), the residue term in (14) becomes close to unity and hence the SWE is compensated for the $k$-th path. It should be noted here that the power for the $k$-th target is enhanced significantly compared to other targets due to near-perfect SWE compensation, and the IF signal corresponding to this target can be approximately recovered by putting an appropriate threshold, i.e., update the $\tilde{y}_{q}[n]$
$\tilde{y}_{q}[n]\leftarrow\begin{cases}\tilde{y}_{q}[n],&\text{if }|\tilde{y}_{q}[n]|\geq\gamma\\ 0,&\text{if }|\tilde{y}_{q}[n]|<\gamma.\end{cases}$
Equivalently, the IF signal for the $k$-th path, converted to the narrowband signal model, can be written as
@@
This process can be illustrated with an example, as shown in Fig. 2. We consider two targets, where Target 1 is located at $8.5\,\mathrm{m},\,30^{\degree}]$ and Target 2 at $9.5\,\mathrm{m},\,60^{\degree}]$. The system bandwidth is set to $0.05f_{c}$, with $f_{c}=77$ GHz, and the number of fast-time and spatial samples is N=512,Q=64N=512,~Q=64. The two unit-amplitude targets produce spread responses in the angle-delay domain due to the SWE, as depicted in Fig. 2(a). The coarse angle estimates for Target 1 and Target 2 are $30.5^{\degree}$ and $61.045^{\degree}$, respectively. Spatial wideband compensation, performed using (14), is shown in Fig. 2(b) and Fig. 2(c). It can be clearly observed that the power concentration of the respective target in the angle-delay domain significantly improves, and the equivalent narrowband target model, as represented by (15), is effectively achieved.
It is important to note that the accuracy of the wideband effect compensation relies heavily on the precision of the coarse angle estimate. At higher angles and larger bandwidths, the spread caused by SWE becomes sufficiently large to induce notable coarse angle bin migration. As a result, the rotation-based array processing technique proposed in [11] becomes ineffective under these conditions, which we demonstrate numerically in Section 4.
@@
where, $\mathbf{Y}^{w}$ is the stacked measurement matrix of the spatial wideband signal in (10), $\mathbf{\Theta}(\Omega_{\theta})\doteq\mathrm{exp}(j2\pi\frac{\alpha}{N}\Omega_{\theta}qn)$ is the wideband phase matrix, $\mathbf{b,c}$ are steering vectors as defined in (17), and $\circ$ denotes the element-wise product.
After obtaining the refined signature of the $k$-th path, $\{\hat{\hat{a}}^{k},\hat{\hat{\Omega}}_{R}^{k},\hat{\hat{\Omega}}_{\theta}^{k}\}$, we remove the corresponding contribution from the IF signal. This operation ensures that the associated peak is eliminated, allowing the algorithm to detect remaining paths in subsequent iterations. Consequently, we update $\check{y}_{q}[n]$ as follows
@@
4 Results
The normalized bandwidth parameter is set to $\alpha=0.05$, and the carrier frequency is fixed at $f_{c}=77\mathrm{GHz}$. A virtual array consisting of $Q=64$ elements is considered. Additionally, the number of fast-time samples per chirp is set to $N=512$, unless otherwise specified. For the chosen XL-MIMO configuration, the corresponding near-field distance is below the minimum simulated target range, ensuring that all targets lie strictly in the far-field regime and validating the planar wavefront assumption used throughout this work. Moreover, numerical exploration reveals that, the two target blocks remain non-overlapping even if they are separated by $(1^{\degree},~0.1~\mathrm{m})$.
@@ 4 Results
We present an example of localizing $K=2$ targets in Fig. 3. In the spatial narrowband case (Fig. 3(a)), the 2D-MUSIC, 2D-Rotation, and the proposed method all accurately localize the two targets in the range-angle domain. However, under the spatial wideband scenario, the narrowband 2D-MUSIC algorithm detects only one target, while the 2D-Rotation method exhibits a localization error of approximately $2.5^{\circ}$ for the second target, as illustrated in Fig. 3(b). In contrast, the proposed method successfully localizes both targets with high accuracy, placing them in $0.01^{\circ}$ proximity to their actual positions in the spatial wideband case.
Figure 3: Target localization at (a) $\alpha=0.01$ (b) $\alpha=0.05$.
Figure 4: RMSE of the estimated parameters at $\alpha=0.05$ (a) range (b) AoA (c) complex coefficient.
Figure 5: (a) Probability of detection and (b) probability of false alarm at $\alpha=0.05$.
@@
We evaluate the effectiveness of our proposed method by measuring its probability of detection, probability of false alarm, and root mean square error (RMSE). The target scenario consists of $K=3$ far-field point targets with ranges uniformly drawn from the interval $8\,\mathrm{m},20\,\mathrm{m}]$ and angles of arrival (AoAs) uniformly drawn from $10^{\circ},80^{\circ}]$. The target complex gains are modeled as $a_{k}=\exp{\left(j\psi_{k}\right)}$, where $\psi_{k}\sim\mathcal{U}[0,2\pi)$, while the noise samples are drawn as $w_{q}[t]\sim\mathcal{CN}(\mathbf{0},\sigma^{2}\mathbf{I})$, independently across all virtual array elements and time samples.
We evaluate the performance of the proposed CS-based estimation method for XL-MIMO FMCW radar under the SWE by comparing it with three benchmark techniques: the rotation-based spatial wideband estimation method from [11], the 2D-MUSIC algorithm [27], and the conventional 2D-OMP method [28]. The comparison is conducted in terms of target detection probability and parameter estimation accuracy.
A detected target is classified as a hit if its estimated range and AoA fall within the predefined resolution limits ($0.02$ m and $1.5^{\circ}$ in this experiment); otherwise, it is considered a false alarm. The probability of detection and false alarm for different methods is presented in Fig. 5, while the corresponding RMSE values, computed for correctly detected targets, are shown in Fig. 4.
As shown in Fig. 4(a), the proposed method achieves substantially lower range RMSE across the entire SNR range, with rapid error reduction beyond 15-15 dB, while 2D-MUSIC and 2D-OMP exhibit consistently higher errors due to their inability to handle the SWE-induced coupling. The 2D-Rotation method improves upon these narrowband techniques but remains inferior to the proposed approach.
@@
Fig. 5 illustrates the target detection performance in terms of probability of detection and false alarm versus SNR for $\alpha=0.0.5$. As shown in Fig. 5(a), the proposed method achieves a probability of detection close to 1 at SNRs above 10-10 dB, with a rapid performance improvement as SNR increases. In contrast, the 2D-Rotation method maintains a moderate probability of detection of approximately 70–80% across the entire SNR range, while the 2D-OMP method performs inferior to both the proposed and the 2D-Rotation method. Furthermore, the 2D-MUSIC consistently fails to reliably detect targets under spatial wideband conditions, achieving a probability of detection below 50%.
In terms of false detection, depicted in Fig. 5(b), the proposed method exhibits a sharp reduction in probability of false alarm with increasing SNR, achieving near-zero false alarms beyond -15 dB. The 2D-Rotation method and 2D-OMP, however, retains a high probability of false alarm of around 20%, while 2D-MUSIC show persistently high probability of false alarm (around 40–60%) across all SNR values. These results clearly demonstrate the superiority of the proposed compressive sensing-based estimation technique for SWE-affected XL-MIMO FMCW radar, delivering both high detection reliability and low probability of false alarm, even under challenging spatial wideband conditions.
@@
Using the 2D-MUSIC algorithm for 2D range-angle parameter estimation with $Q=64$, $N=512$, and $K=3$ takes approximately 54 seconds on a 10th-generation Intel i7 processor. In comparison, the 2D-Rotation method performs the same task in just 0.57 seconds. Both of these methods, however, are limited to spatial narrowband scenarios. The proposed method not only reduces the estimation time to 0.17 seconds but also extends applicability to spatial wideband cases, offering both improved efficiency and broader utility.
5 Conclusion
In this work, we have proposed a low-complexity joint range-angle estimation methodology for XL-MIMO FMCW radars under the SWE, capable of simultaneously estimating the number of targets and their complex reflection coefficients. Numerical results demonstrate that the proposed technique achieves a significantly higher probability of detection and a lower probability of false alarm rate, along with lower MSE, compared to existing methods. Moreover, the scalability and computational efficiency of the proposed approach make it well-suited for a range of localization-based applications across both low and high-bandwidth settings. Additionally, the proposed algorithm is computationally efficient, as it does not require any supplementary mechanisms to mitigate range-angle bin migration effects induced by SWE.
Future work will focus on extending the proposed framework to jointly address the spatial wideband and near-field effects, enabling accurate parameter estimation for the XL-MIMO radar operating in short-range regimes. Additionally, the current assumption of non-overlapping SWE-induced target blocks will be relaxed by developing robust algorithms capable of resolving overlapping scatter responses.